OpenStudy (anonymous):

If three coins are flipped, what is the probability that there are more heads than tails?

OpenStudy (anonymous):

3/6

OpenStudy (anonymous):

1/2

OpenStudy (anonymous):

three heads and three tails

OpenStudy (anonymous):

3/6

OpenStudy (anonymous):

1/2

OpenStudy (anonymous):

that is the answer

OpenStudy (anonymous):

six flips can either end up being 3 heads or 3 tails

OpenStudy (anonymous):

3/6

OpenStudy (anonymous):

1/2

OpenStudy (anonymous):

that is the answer, right?

OpenStudy (anonymous):

2 can be a multiple of itself, right?

OpenStudy (chihiroasleaf):

3/8 ?

OpenStudy (anonymous):

1/2

OpenStudy (anonymous):

3/6

OpenStudy (anonymous):

3 heads and 3 tails

OpenStudy (anonymous):

3/6

OpenStudy (anonymous):

1/2

OpenStudy (anonymous):

all evens = 2^3 = 8

OpenStudy (anonymous):

there is a 50 percent chance of getting more heads than tails, right?

OpenStudy (anonymous):

The chances are always 1/2 so.. 1/2 x 1/2 x 1/2 = 1/8 ???

OpenStudy (anonymous):

are am I thinking of it the wrong way?

OpenStudy (anonymous):

3 heads and 3 tails

OpenStudy (anonymous):

3/6

OpenStudy (anonymous):

right?

OpenStudy (chihiroasleaf):

this is the sample {TTT, TTH, THT, THH, HTT, HTH, HHT, HHH} ahh..., I think I did a mistake, the events that head is more than tails are THH, HTH, HHT, and HHH so its 4/8 = 1/2

OpenStudy (anonymous):

Is your pic a voodoo doll?

OpenStudy (anonymous):

i think the means there are more heads than tails are "2 heads, 1 else (under)"

OpenStudy (chihiroasleaf):

what is voodoo doll @jiji501 ?

OpenStudy (anonymous):

Look it up ;)

OpenStudy (anonymous):

I understand what you are trying to say

OpenStudy (anonymous):

ty

OpenStudy (chihiroasleaf):

can't we include the all heads in it? @tanjung ?

OpenStudy (anonymous):

Yes, we can.

OpenStudy (anonymous):

oh yea, so it's probability is 4/8 = 1/2